Operations Report – May 13th

Crew 299 Operations Report 05-13-2024

SOL: 1

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue), Observatory Lower Deck (broken)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission. Issue with the Observatory Lower Deck was noted before our mission, will be manually opened and closed.

No Rovers used.
General notes on rovers: NA

Summary of Hab operations:
WATER USE: NA, first day of measurements
Water (static tank): 530 (estimated) first day of measurement
Static tank pipe heater: off
Static tank heater: off
Toilet tank emptied: yes

Summary of Internet usage: Catch up on Earthly news, email responses, code learning + development

Summary of suits and radios:
Radios 1,2,3,4,5 and 6 were used
EVA Suits 7,8, and 10 were used (worn for suit adjustments and taken off)

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: Soil exfoliation and checking on radishes

Summary of RAM operations:
The astronomy laptop had a failed boot drive, the drive was extracted and successfully booted on another Windows machine. All software from the Astronomy laptop is usable on a new interface.
Fixed missing zip ties: 4 replaced on the tunnel to the RAM, 4 replaced on the main tunnel, 3 replaced on the tunnel to ScienceDome, 3 replaced on the tunnel to Observatory
Tools Used: Multimeter (suit power check), Measuring tape: Initial water level check, Scissors to cut zip ties

Summary of observatory issues: Temporary issue with the upper deck failing to close as a result of a broken lower deck. Needed to manually press down the lower shutter limit switch in order to trigger the closing motor and mechanism for the upper deck.

Copyright © The Mars Society. All rights reserved. | Main Site